40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Dec 4 02:59:05 2014, ericq, Update, LSC, Bump in Darm Plant Dec4_darmbump.pdf
    Reply  Thu Dec 4 03:09:06 2014, Jenne, Update, LSC, Bump in Darm Plant: Lockloss after measurement 8x
Message ID: 10755     Entry time: Thu Dec 4 03:09:06 2014     In reply to: 10754
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Bump in Darm Plant: Lockloss after measurement 

We were sitting around arm powers of 6, and that loop measurement had finished.  I was about to go down to arm powers of 5ish, but we lost lock.  I'm not sure why.  There's some slow stuff going on in some of the servos, but nothing jumps out at me as a loop oscillation.  It does however kind of look like the PRMI lost lock just before the arm powers went down?  Perhaps this somehow triggered a lockloss?

The time is 1101721675.

Wide view plots:

Arms6_3Dec2014_powers.png

Arms6_3Dec2014_errctrl.png

Arms6_3Dec2014_auxerr.png

Arms6_3Dec2014_angles.png

Zooms:

Arms6_3Dec2014_powers_Zoom.png

Arms6_3Dec2014_errctrl_Zoom.png

Arms6_3Dec2014_auxerr_Zoom.png

Arms6_3Dec2014_angles_Zoom.png


We're stopping for tonight because ETMX is back to its lame-o jumping around.  I went in and squished the cables, but it's still happening.

Also, the FSS PC drive has been high the last few minutes (only starting after we quit for the night).  When the MC re-locks, it sounds like an ocean wave dying out as the noise goes down a little bit.  But, after a few minutes, it'll get mad again and unlock the MC.

Also, also, I noticed this on Monday with Diego, but the LSC-ALS[x,y] filter module gains sometimes mysteriously get set to zero.  WTF?  Eric and I have both independently checked, and we cannot find a single script in the scripts directory with the string "LSC-ALS", so we aren't deliberately changing those.  Does anyone know what might be going on here?

ELOG V3.1.3-